# Channel-Wise Mixed-Precision Quantization for Large Language Models

### v2

@@ 4.1 Experiment Setup
LLM Models and Datasets. We perform our experiments on two models from the OPT family (Zhang et al., 2022) (OPT-2.7B and OPT-6.7B) and two models from the LLaMA2 family (Touvron et al., 2023b) (LLaMA2-7B and LLaMA2-13B). The evaluation of the quantized models is based on perplexity across two language generation tasks, WikiText-2 (Merity et al., 2016) and C4 (Raffel et al., 2020), as perplexity is a widely recognized metric for assessing the LLM performance. For calibration, we follow previous works (Chee et al., 2024; Frantar et al., 2022) and use a set of 128 randomly selected 2048-token segments from the C4 dataset, which contains generic text data from web crawls. This ensures consistency in comparison with baselines and avoids the use of task-specific data when quantizing other datasets. All experiments are implemented in PyTorch (Paszke et al., 2019) and executed on two A6000 GPUs, with performance monitoring handled by the Torch CUDA profiler. we extend our evaluation in Appendix A.3 to include quantization results for other OPT models, scaling up to 30B parameters, as well as a newer model, LLaMA3-8B (AI@Meta, 2024).
@@ 4.4 Comparison with SqueezeLLM
As shown in Table 3, SqueezeLLM outperforms CMPQ across various models at different quantization levels. However, this improvement comes at a significant cost: SqueezeLLM requires four times the memory for the quantization process, making it impractical for larger models, especially under resource constraints. In contrast, CMPQ offers a more efficient solution when memory is limited, requiring only 1/4 of the computational resources. Moreover, if an additional 10% of storage space is available, CMPQ can achieve better performance than SqueezeLLM, particularly in low-memory environments. At higher precision, such as 4-bit quantization, the tradeoff between computational resource requirements and model storage is less pronounced. The maximum performance gain of SqueezeLLM over CMPQ is marginal – only $(5.61-5.57)/5.61=0.71\%$. This minimal improvement renders the substantial additional resource demands of SqueezeLLM unnecessary.
Overall, while CMPQ may not surpass SqueezeLLM in the integer-only quantization tasks, the method offers significant advantages by eliminating the 300% increase in computational overhead associated with SqueezeLLM’s gradient-based approach, with only a 10% increase in storage. Furthermore, CMPQ is more versatile, as it applies to non-integer bit quantization tasks, making it a more practical option for a wider range of scenarios.
@@ 4.6 Ablation Study of Outlier Protection
5 Conclusion and Future Work
In this work, we focused on mixed-precision quantization and aimed to design an algorithm capable of adapting to any bit-width constraint. We observed that different weight channels had varying impacts on model performance, and that activation distributions helped identify salient channels. Building on these insights, we proposed CMPQ, which integrated a channel-wise non-uniform quantization strategy. To further enhance performance, CMPQ introduced two types of outliers that collaboratively preserved critical information. Experimental results showed that CMPQ harnessed the potential of mixed-precision quantization in two key ways: (1) it achieved superior performance in integer-bit quantization tasks, and (2) it delivered significant performance improvements with only a modest increase in memory requirements. In the future, our focus will be on deploying CMPQ-quantized LLMs on real-world devices. This will involve addressing key challenges such as engineering hardware acceleration and designing efficient lookup table kernels to optimize performance.

### v3

@@ 4.1 Experiment Setup
LLM Models and Datasets. We perform our experiments on two models from the OPT family (Zhang et al., 2022) (OPT-2.7B and OPT-6.7B) and two models from the LLaMA2 family (Touvron et al., 2023b) (LLaMA2-7B and LLaMA2-13B). The evaluation of the quantized models is based on perplexity across two language generation tasks, WikiText-2 (Merity et al., 2016) and C4 (Raffel et al., 2020), as perplexity is a widely recognized metric for assessing the LLM performance. For calibration, we follow previous works (Chee et al., 2024; Frantar et al., 2022) and use a set of 128 randomly selected 2048-token segments from the C4 dataset, which contains generic text data from web crawls. This ensures consistency in comparison with baselines and avoids the use of task-specific data when quantizing other datasets. All experiments are implemented in PyTorch (Paszke et al., 2019) and executed on two A6000 GPUs, with performance monitoring handled by the Torch CUDA profiler. we extend our evaluation in Appendix A.3 to include quantization results for other OPT models, scaling up to 30B parameters, as well as a newer model, LLaMA3-8B (AI@Meta, 2024). We further evaluate the domain-specific knowledge and problem-solving ability using MMLU (Hendrycks et al., 2021) in Appendix A.3.
@@
Latency Profiling. We measure the latency and peak memory usage for generating 128 and 1024 tokens on an A6000 GPU using the Torch CUDA profiler. Following Kim et al. (2024), we implement a standard kernel for single-batch inference based on the widely used open-source codebase GPTQ-For-LLaMA. Details of the implementation and comparison can be found in Appendix B.
@@ 4.4 Comparison with SqueezeLLM
As shown in Table 3, SqueezeLLM outperforms CMPQ across various models at different quantization levels. However, this improvement comes at a significant cost: SqueezeLLM requires four times the memory for the quantization process, making it impractical for larger models, especially under resource constraints. In contrast, CMPQ offers a more efficient solution when memory is limited, requiring only 1/4 of the computational resources. Moreover, if an additional 10% of storage space is available, CMPQ can achieve better performance than SqueezeLLM, particularly in low-memory environments. Overall, while CMPQ may not surpass SqueezeLLM in the integer-only quantization tasks, the method offers significant advantages by eliminating the 300% increase in computational overhead associated with SqueezeLLM’s gradient-based approach, with only a 10% increase in storage. Furthermore, CMPQ is more versatile, as it applies to non-integer bit quantization tasks, making it a more practical option for a wider range of scenarios.
@@ 4.6 Ablation Study of Outlier Protection
5 Conclusion
In this work, we focused on mixed-precision quantization and aimed to design an algorithm capable of adapting to any bit-width constraint. We observed that different weight channels had varying impacts on model performance, and that activation distributions helped identify salient channels. Building on these insights, we proposed CMPQ, which integrated a channel-wise non-uniform quantization strategy. To further enhance performance, CMPQ introduced two types of outliers that collaboratively preserved critical information. Experimental results showed that CMPQ harnessed the potential of mixed-precision quantization in two key ways: (1) it achieved superior performance in integer-bit quantization tasks, and (2) it delivered significant performance improvements with only a modest increase in memory requirements. Besides, our latency analysis shows that CMPQ achieves comparable latency and memory usage to the integer-based quantization method.
@@ A.3 Additional Experimental Results
Experimental Results on MMLU Benchmark.
To further demonstrate the effectiveness of CMPQ in complex tasks, we have conducted additional experiments of LLaMA2-7B on the MMLU benchmark, which evaluates a broader range of real-world capabilities. In particular, we focus on in-context learning scenarios, i.e., the 5-shot performance on the MMLU benchmark (Hendrycks et al., 2021). The results, presented in Table 9, illustrate that CMPQ consistently achieves higher accuracy across various bit-widths compared to all baselines. This demonstrates that CMPQ not only excels in perplexity evaluation but also performs robustly in more challenging scenarios.
Table 9: Comparison of PTQ methods on 5-shot MMLU accuracy applied to LLaMA2-7B.
Method Avg. Bit STEM Social Humanities Other Average
FP16 16 37.31 52.91 43.04 53.82 46.46
RTN 2 23.82 23.43 24.52 26.8 24.65
GPTQ 2 23.0 23.9 24.3 26.1 24.2
AWQ 2
QuIP 2
CMPQ 2 25.35 23.04 25.74 26.25 25.18
RTN 3 27.93 23.63 24.87 24.12 25.08
GPTQ 3 28.24 25.67 31.69 32.92 29.85
AWQ 3 31.74 36.59 30.86 38.09 33.98
QuIP 3 27.27 30.81 27.08 25.51 27.57
CMPQ 3 32.54 43.26 36.34 45.99 39.27
RTN 4 34.59 46.05 39.62 47.78 41.83
GPTQ 4 34.13 43.87 38.13 45.59 40.25
AWQ 4 36.24 50.31 41.72 52.22 44.85
QuIP 4 36.25 48.07 39.83 49.07 43
CMPQ 4 36.48 50.57 41.91 52.16 45.00
Appendix B Latency and Peak Memory Usage of CMPQ
We measure the latency and peak memory usage for generating 128 and 1024 tokens on an A6000 GPU using the Torch CUDA profiler. For single-batch inference, we adopt the standard kernel implementation from SqueezeLLM (Kim et al., 2024), based on the widely used open-source codebase GPTQ-For-LLaMA. The quantization process is illustrated in Figure 5. Inspired by LLM.int8() (Dettmers et al., 2022), we handle the multiplication of matrices with different bit-widths separately. Outliers in the weight matrix are stored as CSR sparse matrices, with computations performed using a standard CSR-based kernel. Latency comparisons with SqueezeLLM were conducted on LLaMA2 models using 3-bit quantization, with the results presented in Table 10. Despite incorporating a mixed-precision scheme, CMPQ achieves latency and memory usage comparable to the integer-based quantization method used in SqueezeLLM. We also observe that a small number of FP16 outliers contribute only marginally to the overall computational complexity or performance. Furthermore, their impact diminishes as the size of the language model increases. We attribute this to the fact that the primary memory cost during inference is driven by the mixed-precision model itself, rather than the outliers. While CMPQ’s primary objective is to fully utilize moderate storage space to deliver significant performance improvements over integer quantization, it also imposes only minimal deployment requirements.
Figure 5: Schematic of real quantization process of CMPQ. Given FP16 inputs and quantized weights, the features and weights are decomposed into sub-matrices based on the channel-wise precision allocation returned by Algorithm 1. Then each pair of feature and weight sub-matrices are multiplied together. The outlier matrices are multiplied in FP16 with the original input. Finally, both outlier and regular outputs are accumulated in 16-bit floating point outputs.
Table 10: Hardware profiling of latency and memory usage for LLaMA2-7B and LLaMA2-13B quantized to an average of 3 bits, during the generation of 128 tokens on an A6000 GPU. Results are compared with SqueezeLLM using the standard transformer kernel implementation.
Method Latency (Seconds) Peak Memory (GB)
7b 13B 7b 13B
SqueezeLLM 3.9 6.2 3.2 5.8
CMPQ w/o Outlier 3.84 6.84 3.58 6.44
CMPQ 3.96 6.91 3.59 6.44